OpenStudy (ksaimouli):

partial derivative of m sin(2mx+2my)

OpenStudy (ksaimouli):

last part is +2ny

OpenStudy (ipwnbunnies):

This is easier than you think then. I was trying to find the partial derivative symbol, but I don't know where it is lol.

OpenStudy (ksaimouli):

instead use d i will understand

OpenStudy (ipwnbunnies):

So since we're trying to find partial d wrt m, we treat x, m, and y as constants. Then, it becomes a simple calculus 1 problem.

OpenStudy (ipwnbunnies):

dm/d[f(x,y,m,n)] = 2x cos (2mx + 2ny)

OpenStudy (ipwnbunnies):

It's not sigma, it's its own special character, and I'm not sure why lol. Can someone check my derivative?

OpenStudy (ksaimouli):

\[m \sin(2mx+2ny)\]

OpenStudy (ipwnbunnies):

No, it'll be \[2xsin(2mx+2ny)\]

OpenStudy (ksaimouli):

thats the question

OpenStudy (anonymous):

\[ \frac{\partial }{\partial m} \sin(2mx+2my) = \left[\frac{\partial }{\partial (2mx+2my)} \sin(2mx+2my) \right]\left[ \frac{\partial }{\partial m}2mx+2my\right] \]

OpenStudy (ipwnbunnies):

We're treating x, n, and y's as constants in partial derivatives. So, we'll have to use a u-sub on the inside of the argument. u = 2mx + 2ny du/dm = 2x + 0. The term '2ny' is treated as a constant, and we use the Power Rule on the 2xm.

OpenStudy (anonymous):

First step is chain rule.

OpenStudy (ipwnbunnies):

Hmm, I'm not even good at explaining partial derivaives. I'll let wio take over lol.

OpenStudy (ksaimouli):

I used the chain rule\[f_{x}=(m) \cos(2mx+2ny)(2m)\]

OpenStudy (anonymous):

We are doing partial derivative with respect to \(x\)?

OpenStudy (ksaimouli):

yes

OpenStudy (ipwnbunnies):

Oh I thought you said m.

OpenStudy (ipwnbunnies):

Your original post says 'm'.

OpenStudy (anonymous):

Your answer looks ok then.

OpenStudy (ksaimouli):

this actually fxx (i already found fx) that was the part of the question

OpenStudy (ipwnbunnies):

Right. \[f_{x} = 2m*\cos(2xm + 2ny)\]

OpenStudy (ksaimouli):

okay let me give the question \[(\sin(mx+ny))^2\]

OpenStudy (ksaimouli):

so I found fx and fy to be the same m sin(2mx +2ny) and then found fxx (above)

OpenStudy (ksaimouli):

now fyy= same as fxx right?

OpenStudy (anonymous):

Well, it's similar by symmetry.

OpenStudy (ksaimouli):

okay i am stuck at fxy

OpenStudy (anonymous):

do x, then do y

OpenStudy (ksaimouli):

hold on the Q is \[f_x=m \sin(2mx+2ny)\] so \[f_xx= m \cos(2mx+2ny)(2m)+\cos(2mx+2ny)\]

OpenStudy (ksaimouli):

right

OpenStudy (ksaimouli):

@wio

OpenStudy (anonymous):

actually, now that I think about it... you shouldn't have \(2mx+2ny\), there is no reason for the \(2\).

OpenStudy (ksaimouli):

well question is \[\sin^2(mx+ny)\] \[fx=2(\sin(mx+my))(\cos(mx+my)(m)\] \[msin(2mx+2my)\]

OpenStudy (ksaimouli):

sin2u=sinu cosu

OpenStudy (anonymous):

So you used double angle formula?